A force, F1, does W1=3215.0J of work, in 15.0s. A second force, F2 does W1=420.0Jof work, in 2.0s. Which force delivers greater power?


Hint:P=Wt

A) F2;

B) F1;

C) they both deliver the same amount of power;
Answer #1

the relation between power, work and time

Power = work/time

Power =3215/15=214 W

Power = 420/2 = 210
force F1 gives the greater power